# OctreeOcc: Efficient and Multi-Granularity Occupancy Prediction Using Octree Queries

### 09 Dec 2023 v2

@@ 3.5 Octree Encoder
Given octree queries $Q_{octree}$ and extracted image features $\mathbb{F}$, the octree encoder updates both the octree query features and the octree structure.
@@
Spatial-Tenporal Attention. Referring to the querying paradigm in dense query-based methods[41, 33], we adopt efficient deformable attention[48] for image cross-attention(ICA) and temporal self-attention(TSA).
ICA is devised to enhance the interaction between multi-scale image features and octree queries. Specifically, for an octree query $q$, we can obtain its centre’s 3D coordinate $(x,y,z)$ as reference point $Ref_{x,y,z}$. Then we project the 3d point to images like formula 1 and perform deformable attention:
$ICA(q,\textbf{F})=\frac{1}{N}\sum_{n\in N}\sum_{m=1}^{M_{1}}DA(q,\pi_{n}(Ref_{x,y,z}),\textbf{F}_{n})$ (2)
where $N$ denotes the camera view, $m$ indexes the reference points , and $M_{1}$ is the total number of sampling points for each query. $\textbf{F}_{n}$ is the image features of the $n$-th camera view. DA represents deformable attention.
@@
In accurately representing the driving scene, temporal information plays a crucial role. Given the history octree queries $Q_{t-1}$, we align it to the current octree queries $Q_{t}$ by the ego-vehicle motion. To mitigate computational costs, we employ similar operation to ICA. Given octree queries $Q_{t}\in\mathbb{R}^{N,C}$, a single query with feature $q\in\mathbb{R}^{C}$ and the cooresponding coordinate $p=(x,y,z)$, the TSA is represented by:
$TSA(q,p,Q_{t})=\sum_{k=1}^{K}W_{k}\sum_{m=1}^{M_{2}}A_{k,m}W_{m}Q_{t}(p+\Delta p_{km})$ (3)
where $K$ represents the number of attention heads, $M_{2}$ indicates the number of sampling points, $W_{k}$ and $W_{m}$ are the learning weights, $A_{km}$ denotes the normalized attention weight, and $p+\Delta p_{km}$ signifies the learnable sample point position in 3D space. The feature is computed through trilinear interpolation from the voxel feature at this position.
@@
Iterative Structure Rectification Module. While our initial octree structure is derived from information obtained through image segmentation, it may not precisely match the scene due to segmentation errors and projection inaccuracies. Simultaneously, the encoded octree query acquires valuable spatial information and the octree structure predicted from it complements and corrects the initial prediction, enabling us to address the shortcomings resulting from segmentation challenges.
Specifically, we partition the previous octree structure into two parts: the high-confidence region and the low-confidence region. By sorting the octree split probability values stored in the octree mask from largest to smallest and selecting the top K percent of the regions, we identify the location of high-confidence regions. These are areas where we believe the predictions are accurate and do not necessitate additional adjustments to prevent degradation.
@@
For regions where confidence remains low, it is crucial to refine their outcomes by incorporating additional information sources. Given that predicting the octree structure requires a dense query format, we initially transform the octree query into a dense query, denoted as $Q_{dense}\in\mathbb{R}^{X\times Y\times Z\times C}$, using the procedure outlined in Section 3.3. Subsequently, we downsample it to each level of the octree through trilinear interpolation, resulting in $Q_{dense}^{l}\in\mathbb{R}^{\frac{X}{2^{L-l}}\times\frac{Y}{2^{L-l}}\times\frac{Z}{2^{L-l}}\times C}$. By utilizing the index of low-confidence regions, we can extract the corresponding features $Q_{lcr}^{l}\in\mathbb{R}^{N_{l}\times C},l=0,1,\ldots,L-1$. Employing a Multilayer Perceptron (MLP) to predict octree split probabilities from $Q_{lcr}^{l}$, we apply a weighted sum with the previous predictions of low-confidence regions to obtain refined predictions. These refined predictions are then concatenated with the preserved predictions of high-confidence areas, resulting in the final rectified octree mask. The mask facilitates the transformation of the dense query into a rectified octree query for the subsequent layer.
@@ 4 Experiments
Table 1: 3D Occupancy prediction performance on the Occ3D-nuScenes $val$ set. “normal-$\star$” denotes training with the camera mask.
@@
MonoScene[3] EfficientNetB7 CVPR 2022 6.06 1.75 7.23 4.26 4.93 9.38 5.67 3.98 3.01 5.90 4.45 7.17 14.91 6.32 7.92 7.43 1.01 7.65
BEVDet[8] ResNet101 arxiv 2021 11.73 2.09 15.29 0.0 4.18 12.97 1.35 0.0 0.43 0.13 6.59 6.66 52.72 19.04 26.45 21.78 14.51 15.26
BEVFormer[18] ResNet101 ECCV 2022 23.67 5.03 38.79 9.98 34.41 41.09 13.24 16.50 18.15 17.83 18.66 27.70 48.95 27.73 29.08 25.38 15.41 14.46
BEVStereo[15] ResNet101 AAAI 2023 24.51 5.73 38.41 7.88 38.70 41.20 17.56 17.33 14.69 10.31 16.84 29.62 54.08 28.92 32.68 26.54 18.74 17.49
TPVFormer[9] ResNet101 CVPR 2023 28.34 6.67 39.20 14.24 41.54 46.98 19.21 22.64 17.87 14.54 30.20 35.51 56.18 33.65 35.69 31.61 19.97 16.12
OccFormer[45] ResNet101 ICCV 2023 21.93 5.94 30.29 12.32 34.40 39.17 14.44 16.45 17.22 9.27 13.90 26.36 50.99 30.96 34.66 22.73 6.76 6.97
CTF-Occ[35] ResNet101 arxiv 2023 28.53 8.09 39.33 20.56 38.29 42.24 16.93 24.52 22.72 21.05 22.98 31.11 53.33 33.84 37.98 33.23 20.79 18.00
RenderOcc[27] ResNet101 arxiv 2023 26.11 4.84 31.72 10.72 27.67 26.45 13.87 18.2 17.67 17.84 21.19 23.25 63.20 36.42 46.21 44.26 19.58 20.72
${}^{\star}$ Swin-B arxiv 2022 42.02 12.15 49.63 25.1 52.02 54.46 27.87 27.99 28.94 27.23 36.43 42.22 82.31 43.29 54.46 57.9 48.61 43.55
${}^{\star}$ ResNet101 arxiv 2023 42.13 11.67 50.48 29.64 49.44 55.52 23.29 33.26 30.55 30.99 34.43 42.57 83.31 44.23 54.40 56.04 45.94 40.40
${}^{\star}$ ResNet101 ICCV 2023 43.41 12.10 50.23 32.31 48.55 52.89 31.20 31.25 30.78 32.33 37.06 40.22 83.34 49.27 57.13 59.88 47.67 41.76
${}^{\star}$ ResNet101 - 44.02 11.96 51.70 29.93 53.52 56.77 30.83 33.17 30.65 29.99 37.76 43.87 83.17 44.52 55.45 58.86 49.52 46.33
Table 2: Comparison of query form and efficiency on the Occ3D-nuScenes $val$ set. The experiments were conducted on the NVIDIA A100 GPU.
Methods Query Form Query Resolution mIoU Latency Memory
BEVFormer[18] 2D BEV 200$\times$200 23.67 302 ms 25100M
TPVFormer[9] 2D tri-plane 200$\times$(200+16+16) 28.34 341 ms 29000M
PanoOcc[41] 3D voxel 100$\times$100 $\times$16 42.13 502 ms 35000M
FBOCC[19] 3D voxel & 2D BEV 200$\times$200$\times$16 & 200$\times$200 43.41 463 ms 31000M
Ours Octree Query - 44.02 386 ms 26500M
In this section, we commence by introducing the datasets, evaluation metrics, and implementation details. Following this, we present the results and analysis from comprehensive comparison experiments and ablation studies to validate the efficacy and superiority of our method.
4.1 Datasets
Occ3D-nuScenes[35] constitutes a re-annotation of the nuScenes dataset[2], where precise ground truth labels for occupancy have been derived by aggregating LiDAR scans and human annotations. These annotations adhere to the ego coordinate system framework, consisting of 700 instances for training and 150 for validation. The spatial extents of occupancy along the X and Y axes in ego-coordinates are limited to the range of -40 meters to 40 meters, while the Z-axis spans from -1 meter to 5.4 meters. The occupancy labels are categorized into 17 distinct classes, with each class representing a volumetric space of 0.4 meters in each dimension (0.4m × 0.4m × 0.4m). Additionally, the dataset includes semantic labeling for the 18th category, denoted as “free”, which represents the empty regions within a scene. Finally, Occ3D-nuScenes provides visibility masks for both LIDAR and camera modalities.
4.2 Evaluation metrics
4.3 Implementation Details
Based on previous research, we set the input image size to 900$\times$1600 and employ ResNet101-DCN[5] as the image backbone. We extract multi-scale features from the Feature Pyramid Network (FPN)[20] with downsample sizes of 1/8, 1/16, 1/32, and 1/64. The feature dimension $C$ is set to 256. The octree depth is 3, and the initial query resolution is 50$\times$50$\times$4. For $\alpha$, we designate $\alpha^{1}$ as 0.2 and $\alpha^{2}$ as 0.6. The octree encoder comprises three layers, each composed of Temporal Self-Attention (TSA), Image Cross-Attention (ICA), and Iterative Structure Rectification (ISR) modules. Both $M_{1}$ and $M_{2}$ are set to 4. In TSA, we fuse 4 frames with a time interval of 0.5s. In ISR, the top 10% predictions are considered high-confidence in level 1, and 30% in level 2. The loss weights are uniformly set to 1.0. For optimization, we employ Adam[12] optimizer with a learning rate of 2e-4 and weight decay of 0.01. The batch size is 8, and the model is trained for 24 epochs, consuming 3 days on 8 NVIDIA A100 GPUs.
4.4 Main Results
3D Occupancy Prediction In Tab. 1, we provide comparasion experiments with other state-of-the-art occupancy prediction methodds on Occ3d-nus validation set. The performance of FBOCC[19] relies on open-source code, which we evaluate after ensuring consistency in details (utilizing the same backbone, image resolution, and excluding CBGS[47]) for a fair comparison. Performance for other methods are reported in a series of works[27, 41, 35]. Our approach demonstrates superior performance on mIoU compared to previous methods, particularly excelling in foreground classes such as barriers, cars, and buses, as well as in scene structure classes like manmade and vegetation. This highlights that processing scenes using multiple granularities aligns better with the scene characteristics, enhancing overall expressiveness.
Moreover, we evaluate the efficacy of our approach by comparing it to alternative methods utilizing diverse query forms, including 2D BEV query, 2D tri-plane query, and 3D query, as depicted in Tab 2. The results indicate that our approach not only surpasses these methods in terms of performance but also demonstrates significantly reduced memory usage and lower latency compared to dense queries, approaching the levels observed with 2D queries. Clearly, sparse octree queries effectively mitigate computational overhead while ensuring robust performance.
Table 3: 3D Semantic Scene Completion performance on SemanticKITTI $val$ set.
Method Published IoU mIoU
@@ 4.4 Main Results
LMSCNet[31] 3DV 2020 28.61 6.70 40.68 18.22 4.38 0.00 10.31 18.33 0.00 0.00 0.00 0.00 13.66 0.02 20.54 0.00 0.00 0.00 1.21 0.00 0.00
AICNet[14] CVPR 2020 29.59 8.31 43.55 20.55 11.97 0.07 12.94 14.71 4.53 0.00 0.00 0.00 15.37 2.90 28.71 0.00 0.00 0.00 2.52 0.06 0.00
3DSketch[4] CVPR 2020 33.30 7.50 41.32 21.63 0.00 0.00 14.81 18.59 0.00 0.00 0.00 0.00 19.09 0.00 26.40 0.00 0.00 0.00 0.73 0.00 0.00
JS3C-Net[43] AAAI 2021 38.98 10.31 50.49 23.74 11.94 0.07 15.03 24.65 4.41 0.00 0.00 6.15 18.11 4.33 26.86 0.67 0.27 0.00 3.94 3.77 1.45
MonoScene[3] CVPR 2022 36.86 11.08 56.52 26.72 14.27 0.46 14.09 23.26 6.98 0.61 0.45 1.48 17.89 2.81 29.64 1.86 1.20 0.00 5.84 4.14 2.25
TPVFormer[9] CVPR 2023 35.61 11.36 56.50 25.87 20.60 0.85 13.88 23.81 8.08 0.36 0.05 4.35 16.92 2.26 30.38 0.51 0.89 0.00 5.94 3.14 1.52
VoxFormer[17] CVPR 2023 44.02 12.35 54.76 26.35 15.50 0.70 17.65 25.79 5.63 0.59 0.51 3.77 24.39 5.08 29.96 1.78 3.32 0.00 7.64 7.11 4.18
OccFormer[45] ICCV 2023 36.50 13.46 58.84 26.88 19.61 0.31 14.40 25.09 25.53 0.81 1.19 8.52 19.63 3.93 32.63 2.78 2.82 0.00 5.61 4.26 2.86
Symphonies[10] arxiv 2023 41.44 13.44 55.78 26.77 14.57 0.19 18.76 27.23 15.99 1.44 2.28 9.52 24.50 4.32 28.49 3.19 8.09 0.00 6.18 8.99 5.39
Ours - 44.71 13.12 55.13 26.74 18.68 0.65 18.69 28.07 16.43 0.64 0.71 6.03 25.26 4.89 32.47 2.25 2.57 0.00 4.01 3.72 2.36

### 19 Mar 2024 v3

@@ 3.5 Octree Encoder
Given octree queries $Q_{octree}$ and extracted image features $\mathbb{F}$, the octree encoder updates both the octree query features and the octree structure. Referring to the querying paradigm in dense query-based methods[50, 39], we adopt efficient deformable attention[59] for temporal self-attention(TSA) and image cross-attention(ICA).
@@
In accurately representing the driving scene, temporal information plays a crucial role. By leveraging historical octree queries $Q_{t-1}$, we align it to the current octree queries by the ego-vehicle motion transformation. Given historal octree queries $Q_{t-1}\in\mathbb{R}^{N,C}$, a current octree query $q$ located at $p=(x,y,z)$, the TSA is represented by:
$TSA(q,Q_{t-1})=\sum_{m=1}^{M_{1}}DeformAttn(q,p,Q_{t-1}),$ (2)
where $M_{1}$ indicates the number of sampling points. Implementing it within the voxel-based self-attention ensures that each octree query interacts exclusively with local voxels of interest, keeping the computational cost manageable.
Image cross-attention is devised to enhance the interaction between multi-scale image features and octree queries. Specifically, for an octree query $q$, we can obtain its centre’s 3D coordinate $(x,y,z)$ as reference point $Ref_{x,y,z}$. Then we project the 3D point to images like formula 1 and perform deformable attention:
$ICA(q,\textbf{F})=\frac{1}{N}\sum_{n\in N}\sum_{m=1}^{M_{2}}DeformAttn(q,\pi_{n}(Ref_{x,y,z}),\textbf{F}_{n}),$ (3)
where $N$ denotes the camera view, $m$ indexes the reference points , and $M_{2}$ is the total number of sampling points for each query. $\textbf{F}_{n}$ is the image features of the $n$-th camera view.
Figure 3: Illustration of octree structure rectification. The left figure displays the initially predicted octree structure and the right figure depicts the octree structure after Iterative Structure Rectification. It is evident that the predicted octree structure becomes more consistent with the object’s shape following the rectification module.
@@
Iterative Structure Rectification Module. The initial octree structure is derived from information obtained through image segmentation, but it may not precisely correspond to the actual scene due to potential segmentation and projection errors. Nonetheless, the encoded octree query captures crucial spatial information. Thus, the predicted octree structure based on this query complements and rectifies the initial structure prediction, allowing us to mitigate limitations arising from segmentation and projection issues.
@@
Specifically, we partition the octree structure into two parts: the high-confidence regions and the low-confidence regions, as Fig. 2 shows. By sorting the octree split probability values stored in the octree mask in descending order and selecting the top kof regions at each level, we can identify the locations of high-confidence regions. For these regions, predictions are relatively more accurate and no additional adjustments are required in this iteration. For regions where confidence remains low, we first extract the query features corresponding to those areas by utilizing the index of low-confidence regions, denoted as $Q_{lcr}=\{Q_{lcr}^{l}\in\mathbb{R}^{N_{l}\times C}\}_{l=1}^{L-1}$, where $N_{l}$ represents the number of low-confidence queries in level $l$. We then employ a Multi-layer Perceptron (MLP) to predict octree split probabilities from $Q_{lcr}$. Subsequently, we apply a weighted sum with the previous split probability predictions of low-confidence regions to obtain rectified predictions. These refined predictions are concatenated with the preserved probability predictions of high-confidence regions, culminating in the generation of the final rectified octree mask. More details are shown in Supplementary.
It is worth noting that, due to the iterative nature of structure updates, regions initially considered high confidence may not necessarily remain unchanged, as they might be partitioned into low-confidence regions in the next iteration.
3.6 Octree Ground Truth Generation
We derive the octree ground truth from the semantic occupancy ground truth. Specifically, for a voxel at level $l$ in the octree, we identify its corresponding $8^{L-l}$ voxels in the semantic occupancy ground truth. If these voxels share the same labels, we deem the voxel at level $l$ unnecessary to split (assigned a value of 0); otherwise, it necessitates division (assigned a value of 1), as the current resolution is insufficient to represent it adequately. Through this process, each voxel at each level is assigned a binary value of 0 or 1. Then we obtain the octree ground truth $G_{octree}=\{G_{octree}^{l}\in\mathbb{R}^{\frac{X}{2^{l}},\frac{Y}{2^{l}},\frac{Z}{2^{l}}}\}_{l=1}^{L-1}$. Here, $L$ represents the depth of the octree, while $X$, $Y$, and $Z$ denote the volume resolution of the semantic occupancy ground truth. $G_{octree}$ is employed to supervise the octree mask using focal loss, facilitating the network in learning the octree structure information.
@@ 4 Experiments
In this section, we commence by introducing the datasets (Sec. 4.1), evaluation metrics (Sec. 4.2), and implementation details (Sec. 4.3). Subsequently, we evaluate our method on 3D occupancy prediction and semantic scene completion tasks (Sec. 4.4) to demonstrate its effectiveness. Additionally, we conduct ablation studies (Sec. 4.5) and provide more analysis (Sec. 4.6) of our method to validate its efficacy and superiority.
4.1 Datasets
Occ3D-nuScenes[43] constitutes a re-annotation of the nuScenes dataset[2], where precise ground truth labels for occupancy have been derived by aggregating LiDAR scans and human annotations. These annotations adhere to the ego coordinate system framework, consisting of 700 instances for training and 150 for validation. The spatial extents of occupancy along the X and Y axes in ego-coordinates are limited to the range of -40 meters to 40 meters, while the Z-axis spans from -1 meter to 5.4 meters. The occupancy labels are categorized into 17 distinct classes, with each class representing a volumetric space of 0.4 meters in each dimension (0.4m × 0.4m × 0.4m). Additionally, the dataset includes semantic labeling for the 18th category, denoted as “free", which represents the empty regions within a scene. Finally, Occ3D-nuScenes provides visibility masks for both LIDAR and camera modalities.
4.2 Evaluation metrics
4.3 Implementation Details
Based on previous research, we set the input image size to 900$\times$1600 and employ ResNet101-DCN[5] as the image backbone. Multi-scale features are extracted from the Feature Pyramid Network (FPN)[23] with downsampling sizes of 1/8, 1/16, 1/32, and 1/64. The feature dimension $C$ is set to 256. The octree depth is 3, and the initial query resolution is 50$\times$50$\times$4. We choose query selection ratios of 20% and 60% for the two divisions. The octree encoder comprises three layers, each composed of Temporal Self-Attention (TSA), Image Cross-Attention (ICA), and Iterative Structure Rectification (ISR) modules. Both $M_{1}$ and $M_{2}$ are set to 4. In ISR, the top 10% predictions are considered high-confidence in level 1, and 30% in level 2. The loss weights are uniformly set to 1.0. For optimization, we employ Adam[14] optimizer with a learning rate of 2e-4 and weight decay of 0.01. The batch size is 8, and the model is trained for 24 epochs, consuming around 3 days on 8 NVIDIA A100 GPUs.
Table 1: 3D Occupancy prediction performance on Occ3D-nuScenes dataset. “normal-$\star$” denotes training with the camera mask.
@@ 4.3 Implementation Details
MonoScene[3] ResNet101 CVPR’22 6.06 1.75 7.23 4.26 4.93 9.38 5.67 3.98 3.01 5.90 4.45 7.17 14.91 6.32 7.92 7.43 1.01 7.65
BEVDet[9] ResNet101 arxiv’21 11.73 2.09 15.29 0.0 4.18 12.97 1.35 0.0 0.43 0.13 6.59 6.66 52.72 19.04 26.45 21.78 14.51 15.26
BEVFormer[21] ResNet101 ECCV’22 23.67 5.03 38.79 9.98 34.41 41.09 13.24 16.50 18.15 17.83 18.66 27.70 48.95 27.73 29.08 25.38 15.41 14.46
BEVStereo[19] ResNet101 AAAI’23 24.51 5.73 38.41 7.88 38.70 41.20 17.56 17.33 14.69 10.31 16.84 29.62 54.08 28.92 32.68 26.54 18.74 17.49
TPVFormer[11] ResNet101 CVPR’23 28.34 6.67 39.20 14.24 41.54 46.98 19.21 22.64 17.87 14.54 30.20 35.51 56.18 33.65 35.69 31.61 19.97 16.12
OccFormer[56] ResNet101 ICCV’23 21.93 5.94 30.29 12.32 34.40 39.17 14.44 16.45 17.22 9.27 13.90 26.36 50.99 30.96 34.66 22.73 6.76 6.97
CTF-Occ[43] ResNet101 NeurIPS’23 28.53 8.09 39.33 20.56 38.29 42.24 16.93 24.52 22.72 21.05 22.98 31.11 53.33 33.84 37.98 33.23 20.79 18.00
RenderOcc[32] ResNet101 ICRA’24 26.11 4.84 31.72 10.72 27.67 26.45 13.87 18.2 17.67 17.84 21.19 23.25 63.20 36.42 46.21 44.26 19.58 20.72
${}^{\star}$ Swin-B arxiv’22 42.02 12.15 49.63 25.1 52.02 54.46 27.87 27.99 28.94 27.23 36.43 42.22 82.31 43.29 54.46 57.9 48.61 43.55
${}^{\star}$ ResNet101 CVPR’24 42.13 11.67 50.48 29.64 49.44 55.52 23.29 33.26 30.55 30.99 34.43 42.57 83.31 44.23 54.40 56.04 45.94 40.40
${}^{\star}$ ResNet101 ICCV’23 43.41 12.10 50.23 32.31 48.55 52.89 31.20 31.25 30.78 32.33 37.06 40.22 83.34 49.27 57.13 59.88 47.67 41.76
${}^{\star}$ ResNet101 N/A 44.02 11.96 51.70 29.93 53.52 56.77 30.83 33.17 30.65 29.99 37.76 43.87 83.17 44.52 55.45 58.86 49.52 46.33
Table 2: 3D Semantic Scene Completion performance on SemanticKITTI dataset.
Method Reference IoU mIoU
@@
LMSCNet[36] 3DV’20 28.61 6.70 40.68 18.22 4.38 0.00 10.31 18.33 0.00 0.00 0.00 0.00 13.66 0.02 20.54 0.00 0.00 0.00 1.21 0.00 0.00
AICNet[17] CVPR’20 29.59 8.31 43.55 20.55 11.97 0.07 12.94 14.71 4.53 0.00 0.00 0.00 15.37 2.90 28.71 0.00 0.00 0.00 2.52 0.06 0.00
3DSketch[4] CVPR’20 33.30 7.50 41.32 21.63 0.00 0.00 14.81 18.59 0.00 0.00 0.00 0.00 19.09 0.00 26.40 0.00 0.00 0.00 0.73 0.00 0.00
JS3C-Net[52] AAAI’21 38.98 10.31 50.49 23.74 11.94 0.07 15.03 24.65 4.41 0.00 0.00 6.15 18.11 4.33 26.86 0.67 0.27 0.00 3.94 3.77 1.45
MonoScene[3] CVPR’22 36.86 11.08 56.52 26.72 14.27 0.46 14.09 23.26 6.98 0.61 0.45 1.48 17.89 2.81 29.64 1.86 1.20 0.00 5.84 4.14 2.25
TPVFormer[11] CVPR’23 35.61 11.36 56.50 25.87 20.60 0.85 13.88 23.81 8.08 0.36 0.05 4.35 16.92 2.26 30.38 0.51 0.89 0.00 5.94 3.14 1.52
VoxFormer[18] CVPR’23 44.02 12.35 54.76 26.35 15.50 0.70 17.65 25.79 5.63 0.59 0.51 3.77 24.39 5.08 29.96 1.78 3.32 0.00 7.64 7.11 4.18
OccFormer[56] ICCV’23 36.50 13.46 58.84 26.88 19.61 0.31 14.40 25.09 25.53 0.81 1.19 8.52 19.63 3.93 32.63 2.78 2.82 0.00 5.61 4.26 2.86
Symphonies[12] CVPR’24 41.44 13.44 55.78 26.77 14.57 0.19 18.76 27.23 15.99 1.44 2.28 9.52 24.50 4.32 28.49 3.19 8.09 0.00 6.18 8.99 5.39
Ours N/A 44.71 13.12 55.13 26.74 18.68 0.65 18.69 28.07 16.43 0.64 0.71 6.03 25.26 4.89 32.47 2.25 2.57 0.00 4.01 3.72 2.36
Figure 4: Qualitative results on Occ3D-nuScenes validation set, where the resolution of the voxel predictions is 200$\times$200$\times$16.
4.4 Results
3D Occupancy Prediction. In Tab. 1, we provide comparasion experiments with other state-of-the-art occupancy prediction methodds on Occ3d-nus validation set. The performance of FBOCC[22] relies on open-source code, which we evaluate after ensuring consistency in details (utilizing the same backbone, image resolution, and excluding CBGS[58]) for a fair comparison. Performance for other methods are reported in a series of works[32, 50, 43]. Our approach demonstrates superior performance on mIoU compared to previous methods, particularly excelling in foreground classes such as barriers, cars, and buses, as well as in scene structure classes like manmade and vegetation. This highlights that processing scenes using multiple granularities aligns better with the scene characteristics, enhancing overall expressiveness.
Moreover, we evaluate the efficiency of our approach by comparing it to alternative methods utilizing diverse query forms, as depicted in Tab 3. The results indicate that our approach not only surpasses these methods in terms of performance but also demonstrates significantly reduced memory usage and lower latency compared to dense queries, approaching the levels observed with 2D queries. Clearly, sparse octree queries effectively mitigate computational overhead while ensuring robust performance. Fig. 4 displays qualitative results obtained by our methods and some other methods, illustrating that our approach comprehensively understands the structure of the scene, showcasing superior performance in scene understanding.